OpenStudy (anonymous):

do the equations x =4y + 1 and x =4y -1 have the same solution?

OpenStudy (anonymous):

If you substitute a value into the equations they will not have the same solution, if this is what you are asking. If x=2, the value of y will be 1/4 in the first equation. Whereas in the second equation if x=2, the value if y will be 3/4

OpenStudy (usukidoll):

let y be any number.For example, y=0 x = 4(0)+ 1 = 1 x = 4(0) - 1 = -1 We have established a contradiction.

OpenStudy (campbell_st):

the easy way to look at these linear equations is to write them in slope intercept form by making y the subject 1st equation \[y = \frac{1}{4} x - 1\] 2nd equation, \[y = \frac{1}{4} x + 1\] the lines are parallel, since the both have a slope of 1/4 and different y intersepts, so they have different solutions.